ID |
Date |
Author |
Type |
Category |
Subject |
8968
|
Mon Aug 5 19:10:01 2013 |
Koji | Update | CDS | Choosing LO Amplitudes and Frequencies | - I suppose the green transmission paths were thoroughly inspected and aligned in prior to the measurement
- Of course it is a BAD idea to use 60Hz as the LO frequency.
- Power spectra should be plotted as "RIN (relative intensity noise)" as the DC of 1 and 100 gives you 100 times different power spectra for the same beam.
Don't forget to subtract the offset from your DC values. |
8967
|
Mon Aug 5 18:48:44 2013 |
gautam | Configuration | endtable upgrade | Full range calibration of PZT mounted mirrors for Y-endtable | I had prepared two more PZT mounted mirrors for the Y-end some time back. These are:
- A 2-inch CVI mirror (45 degree, HR and AR for 532nm, was originally one of the steering mirrors at the X-endtable, and was removed while switching those out for the PZT mounted mirrrors).
- A 1-inch Laseroptik mirror (45 degree, HR and AR for 532nm).
I used the same QPD set-up and the methodology described here to do a full-range calibration of these PZTs. Plots attached. The calibration constants have been determined to be:
CVI-pitch: 0.316 mrad/V
CVI-yaw: 0.4018 mrad/V
Laseroptik pitch: 0.2447 mrad/V
Laseroptik yaw: 0.2822 mrad/V
Remarks:
- These PZTs, like their X-end counterparts, showed evidence of drift and hysteresis. We just have to deal with this.
- One of the PZTs (the one on which the CVI mirror is mounted) is a used one. While testing it, I thought that its behaviour was a little anomalous, but the plots do not seem to suggest that anything is amiss.
Plots:
CVI YAW CVI PITCH

Laseroptik YAW Laseroptik PITCH
|
8966
|
Mon Aug 5 18:18:32 2013 |
gautam | Update | CDS | Choosing LO Amplitudes and Frequencies | In order to decide what frequencies to dither the 4 degrees of freedom (M1-pitch&yaw, M2-pitch&yaw) at, I took the power spectrum of the X and Y-arm green transmission (C1:ALS-TRX_OUT, C1:ALS-TRY_OUT). Plots showing the power spectra are attached. Looking at the power spectra, I would think that for the X-arm, it would be okay to dither at 40, 50, 60 and 70 Hz. In order to check if the piezos could respond to these frequencies, I used my QPD setup and shook the PZTs with a 100Hz, 1Vpp sinusoid, and saw that the spot moved smoothly on the QPD.
As for choosing the modulation amplitude, I did a simplistic approximation assuming that the misalignment only rotates the beam axis relative to the cavity axis, and determined what angle coupled 10% of the power into the next eigenmode. Assuming that this is small enough such that if we are already locked to TEM00, the dither won't kick it up to some higher-order mode, the LO amplitude should be in the range of 30-60 digital counts (determined using the PZT calibration constants determined here. This corresponds to a sine-wave of ~50mV amplitude reaching the PZTs (after HV amplification). I am not sure if this is too small, but according to the PZT datasheet, these platforms are supposed to have a resolution of 0.02 urad, which would correspond to the input signal changing by ~0.1 mV, so this signal should be capable of dithering the tip-tilt.
I have already added band-pass filters centered at these frequencies to the model (with a passband of 5Hz, 2Hz on either side), and low-pass filters to pull out the DC component of the output of the lock-in amplifiers. It remains to tune the gains of the filter stages. These parameters (frequency, amplitude of the LOs) may also have to be changed after tests). Hopefully the PZTs can be plugged in tomorrow, and I can try and make a measurement of the output matrix.
Koji also suggested that it may be good to have a path in the model that feeds back to the PZTs by dithering the cavity mirrors as opposed to the PZT mounted mirrors. I will work on incorporating this into the SIMULINK model (c1asx.mdl) and also into the master medm screen.
Notes:
- The spot size of the X-arm green transmission on the PD was larger than the active surface. I moved the GTRX PD a little back and put in a lens (KPX085, 62.9mm FL, AR.14) in front of the PD, such that the spot is now occupying about 1/4th of the active surface area. The lens was mounted in a Thorlabs LMR1mount, and has been labelled.
- I made a slight change to the SIMULINK model, so as to calibrate the PZT sliders to (approximately) volts (I added a multiplier block that multiplies the slider value by constant value 3267.8). The idea is that we can approximately relate the slider value to tilt, knowing the calibration constant in mrad/V for the PZTs.
Power Spectra of Arm Green Transmission:

|
8965
|
Mon Aug 5 18:02:34 2013 |
manasa | Update | General | Close up list checked |
-
Center beam on all AS optics
-
GET CAMERA IMAGES OF EVERYTHING
- We must get images right before closing, right after closing, etc.
Make sure REFL is clear
- dither PRM, see motion on AP tables
Make sure AS is clear
- dither BS/ITM, see motion on AP tables
Using IPANG/POS pick-off mirrors, center beams on:
IPPOS
IPANG (aligned low in pitch)
Check green alignment in the arms and make sure the transmitted green reaches the PSL table.
-
Check all OpLevs centered, in and out of vacuum
[Jenne, Manasa]
IPANG needed to be re-aligned today. Heavy doors are in place and bolts tight (torque 25 & 45).
Steve! We are ready for pump down!
I will check the IFO alignment once again early tomorrow morning before Steve starts pumping down.
|
8964
|
Mon Aug 5 11:53:45 2013 |
Evan | Update | ISS | CTN Servo - Explicit Requirement and Proposed Servo | I goofed on the transfer function requirement by not giving you the plant transfer function, which looks to be about 0.014 V/V, independent of frequency (PSL:1278). This needs to be compensated for in the electronic transfer function. |
8963
|
Mon Aug 5 10:50:48 2013 |
Steve | Update | VAC | RGA background | RGA background at day 12 of this vent . The maglev is pumping on the rga through VM2
|
8962
|
Fri Aug 2 22:51:10 2013 |
Jenne | Update | General | All vent tasks complete, just need oplev check | [Manasa, Koji, Jenne]
We went into the BS and IOO chambers, and aligned the green beams such that they came out of the vacuum chamber. The idea here was to get the beams at the same height, but slightly offset in yaw. This required moving the Periscope on BS table, PBS in front of that periscope, the Periscope on the IOO table, and 2 steering mirrors on the IOO table after the 2nd periscope. The tables were not releveled, although we have aligned the full interferometer to this situation, so we do not want to touch the tables. The MC spot positions are still consistent with those measured earlier this afternoon, before this work, so I'm not concerned.
We confirmed that both green beams are hitting a good place (centered in pitch, and just left and right of center in yaw) on the mirror in the OMC chamber, and are getting to the center of the first mirror on the PSL table. We then coarsely aligned the beams on the PSL table.
We then relocked and aligned the arms for IR, and checked that the AS beam is centered on the mirrors in the BS chamber, and that the beam is coming out, and to the AS table. I touched the last mirror before PZT3 a small amount in yaw, and then PZT3 in pitch and yaw, until we saw the beam recentered on the first mirror on the AS table. At that point, we were also back to the center of the AS camera (which is good, since Koji had aligned all of that the other day). So, the AS beam is good.
We checked IPPOS, and have centered the beam on all the mirrors, and aligned the beam onto the QPD.
We checked IPANG, by looking through the viewports at the mirrors in the ETMY chamber. We are now centered in yaw, but clipping a bit low. This is what we want, since we always end up drifting high during the pump-down.
We see a nice, unclipped REFL beam on the camera.
We see a beam a little high on the POP camera, but Koji looked on the table with a card, and saw the beam....we just need to do minor alignment on the out of vac mirrors.
We checked again that the green TEM00 beams from both arms come to the PSL table.
We are getting POX and POY out, since we are using them to lock and align the arms.
Manasa and Koji recovered one clean allen key from the bottom of the chambers, but one remains, as a sacrifice to the vacuum gods.
I believe that, with the exception of checking the oplevs and taking photos of PR3, and the green steering optics, we have finished all of our vent tasks. We should do a quickie alignment on Monday, check the oplevs, take some photos, and put on the heavy doors. Pumping can start either Monday afternoon or Tuesday morning.
|
8961
|
Fri Aug 2 21:59:36 2013 |
Charles | Update | ISS | Finalized ISS Schematic (hopefully) | Attached is the finalized schematic. The general circuit topology should remain the same from this point forward, although individual component values are subject to change. I will also be adding some more annotations to ensure everything on the board is clear.
In general, I have finally included all of the correct components (i.e. front panel switches are now actually switches and front panel LEDs are now included). I also added an external 'Boost' switch, which can be used to enable or disable the boosts. The motivation for including this switch is that one might want to test functionality of the ISS without using the 'fancy' RMS detection and triggering circuitry. Additionally, one can disable the boosts when all the circuitry is stuffed in order to troubleshoot, so it essentially grants the board some flexibility in its operation.
I am now working on the PCB layout and I should hopefully have that done next week. |
8960
|
Fri Aug 2 17:50:10 2013 |
Jenne | Update | General | PR3 wedge angle adjusted | [Jenne, Manasa, Koji]
Earlier today, we locked and aligned both the X and Y arms.
I then went into the BS chamber, put on the BS' aperture, and put an aperture along the AS path. (We had Michelson fringes, so I centered the aperture around the fringes. I used one of the brass ruler things that we use to center the beam on ITMs and ETMs, on a riser. I put this aperture at the edge of the BS table, after the AS beam is launched toward the OMC chamber. The idea was to replace PR3 such that I could get the beam back through the BS aperture, and the brass ruler aperture, in hopes that we would see arm flashes, and not have to open the ITMY and ETMY heavy doors.)
I set references on the table so that I could put PR3 back in its original position, then removed PR3 from the chamber.
Steve set up a HeNe for me, that we pointed through the optic. The ghost beam was very high, indicating (as expected) that the wedge was not perfectly horizontal.
I took the suspension off of the cage and laid it down, as I have in the past.
I removed the optic from the suspension, to try to figure out which was the fat vs. skinny side. I noticed that there are very faint marks on the actual fat and skinny sides of the optic. (Mpral - for the LaserOptik mirrors, look for the faint lines that are the full width of the barrel, not the placement of the arrow which marks the HR side). I put the optic back in (HR side toward the back, fat side on the left (as you look at the face of the optic), which is consistent with the picture in the Optical Layout page of the Wiki, near the bottom.) the optic holder ring.
I put the suspension back on the cage, and saw that the HeNe's ghost beam was now nearly horizontal relative to the straight-through beam. Excellent. Also, the pitch balancing didn't seem to change noticably, which I determined was within "poking" distance of where we need it to be.
I put PR3 back onto the BS table, and adjusted it around until I got the beam through both the BS aperture, and the one on the AS path. As usual, this took quite a while, but as soon as I got through both of those apertures (really at the same place, not close to being through them, but as close as I could tell by eye - this is what took forever), Koji and Manasa saw flashes in the Yarm! Yay!
Since I had to move PR3 in angle a tiny bit, I reset the references, then dogged down PR3. We still had flashes, this time in both arms, so we closed up the light doors.
We have now locked and aligned both arms in IR after the adjustment of PR3, and see both arms' green at 01 or 02. We are about to start checking the green positioning on the periscopes. We will also need to check the AS path, as well as IPPOS and IPANG before we close up. We see REFL on the camera.
Separately - Manasa remembered that 2 clean things were dropped yesterday - a screw, and an allen key. Since they're both Clean, we're not too worried, although she thinks a long-armed person may be able to reach the allen key. |
8959
|
Thu Aug 1 22:58:45 2013 |
Charles | Update | ISS | CTN Servo - Explicit Requirement and Proposed Servo | In PSL elog 1270, Evan elucidated the explicit requirements for the CTN ISS board. Essentially, the transfer function of the ISS should be something like:
TF_mag = (Unstabilized RIN) / (Calculated RIN Requirement)
I took Evan's data and did exactly this. I then designed a servo (using the general design I proposed here) to meet this requirement with a safety factor of ~10. By safety factor, I mean that if the ISS operates exactly according to theory, it should suppress the noise by a factor of 10 more than what is necessary/set out by the requirement. Below is a plot of the loop gain obtained directly from the requirement (the above expression for TF_mag) and the transfer function of the servo I am proposing.

I don't have the actual schematics attached as I was working with a LISO file and have yet to update the corresponding Altium schematic. The LISO file is attached and I will add the schematics later, although one can reference the second link to find a simple drawing. |
8958
|
Thu Aug 1 22:49:31 2013 |
manasa | Update | General | Green status after PR3 flipping investigated | [EricQ, Koji, Manasa]
We opened the BS chamber to check the status of the green beams. The X green has 3 steering mirrors before they hit periscope1 and the Y green transmits through all the optics giving no way to steer it.
We agreed to start fixing the Y green. The wedge angle of PR3 is steering the transmitted beam away in both pitch and yaw. Since we are restricted only to yaw movement (done by moving the periscope), we want he wedge angle to be oriented in the yaw as well.
Right now, the wedge is oriented at about 20-30 deg off (The mark on the side of the mirror does not indicate the wedge). So we see a pitch as well as yaw misalignment in the transmitted beam. The pitch misalignment is making the beam fall off the mirrors in periscope2.
We have decided to get the wedge angle set right for PR3 and redo the alignment for IR. Once we are aligned for the IR, we will modify the green layout. |
8957
|
Thu Aug 1 21:28:09 2013 |
gautam | Update | CDS | Slow channels set-up in ALS | The following slow channels have been added and are now being recorded by FB.
C1:ALS-X_OVEN_TEMP
C1:ALS-Y_OVEN_TEMP
C1:ALS-BEATX_FREQ
C1:ALS-BEATY_FREQ
Details:
In order to integrate the data collected by the Raspberry-Pi from the Y-end doubling oven temperature controller and also the data from the frequency counter which will be hooked up to monitor the beat frequency, Koji helped me set up some slow EPICS record channels (in ALS as we felt this was most appropriate). The procedure for setting up slow channels was as follows (virtually identical to what is detailed in this elog:
- Add the channel names to the file C0EDCU.ini (path = /opt/rtcds/caltech/c1/chans/daq/C0EDCU.ini).
- Make a database (.db) file so that these channels are actually recorded (path = /cvs/cds/caltech/target/c1aux/als.db).
- Restart framebuilder.
- Verify that the channels indeed exist and can be read and written to using ezcaread and ezcawrite.
I will now integrate these channels into my scripts, and make some simple MEDM screens.
|
8956
|
Thu Aug 1 20:58:56 2013 |
gautam | Update | CDS | New model for endtable PZTs-MEDM Screens setup |
I have made some minor changes to the model, made all the MEDM screens, and linked monitors on these to the appropriate channels. I have borrowed heavily from the C1ASS MEDM screens (particularly for the small filter modules-it was convenient to just copy and paste an existing module, and edit the channel names using EMACS/GEDIT), and have edited these to suit the needs of this servo. Some features:
- The feedback signal (only the output of the servo to the PZTs, plus any contribution from the on-screen sliders, and not including the LO output) is monitored with both a slow (using CDS_EPICS_OUTPUT block from the CDS_PARTS library) and fast channel (using Test Point from the same library). The idea is that it would be useful to know the output to the PZTs such that if coarse adjustment ever needs to be done at the endtable, the PZTs can be restored to the middle of its operating range by means of the sliders.
- Sliders are incorporated into the master screen for adjusting the output to the PZTs. There are text-input fields below the sliders as well, which control the same channel.
- I have removed the 4 remaining excitation points to the DAC set up in C1SCX, and have relocated them to channels 12-15 of the DAC in C1ASX.
I think I am now ready to take some measurements and try and optimize this servo. There is no green transmission at the PSL table at the moment, so not much can be done, though the first step would be to take the power spectrum of the error signal, which would help me decide the appropriate frequencies for the LOs. I would then have to add the appropriate filters to the model. The last, and most difficult step, would be the measurement of the output matrix, though Koji has given me some ideas about how this measurement can be done. I also have a template script ready, though I will only finalise this after optimising the servo and running it a couple of times manually.
Attached are screenshots of the MEDM screens.

|
8955
|
Thu Aug 1 18:55:20 2013 |
Alex Cole | Update | Electronics | Preliminary Photodetector Frequency Reponse Measurements |
Quote: |
[Eric, Alex]
We used our setup from yesterday (elog #8940) to measure transimpedance measurements for AS55, REFL11, REFL33, and REFL55, using our Newport 1611 FC-AC as reference. We connected the fibers to their respective telescopes such that the beams focused on their photodetectors, using a multimeter to maximize photodetector DC output. Plots are attached. At first glance, the poles seem to be where they're supposed to be.
Note that the procedure used today is similar to what the eventual automated procedure will be. The main differences are (1) The RF Switch will be used rather than manual switching (2) NWAG4395A will be used to collect data rather than netgpibdata (3) Data will be fit using vectfit4.m and compared to some canonical set.
|
[Alex, Eric]
Today I spent some time mounting the launcher and performing the same data collection for POX11. I think I still need to focus the launcher so the photodetector gets a good signal, but the data from today wasn't too bad. Additionally, I worked on matlab scripts to improve PDFR data analysis.
This time I collected data from the network analyzer using NWAG4395A in the netgpibdata directory. The advantage of this is that the computer tells the network analyzer to perform the sweep as well as retrieving the data.
For analysis, I improved my implementation of vectfit4.m so that it focuses in on the particular photodetector's predicted peaks and thus ignores much of the noise, giving a better fit. The raw data is the red circles in the 2nd attachment, while the fit is the blue line. I also had the program return the frequency value of the peak. For POX11, this was 1.106e+07 Hz.
I also finagled copies of existing programs to enable one to plot multiple transfer functions on the same axes. This function is /users/alex.cole/plottwo.m. I will eventually use this to compare new data to some canonical data so that we may monitor photodetector performance over time.
The eventual plan is to generate two plots per photodetector, one of which will compare new data to the canonical set, the other of which will show the fit of the data. Both will have subplots that zoom in around regions of interest (known peaks and notches), and the plot which displays the canonical set will also have Q's of peaks and their locations. |
8954
|
Thu Aug 1 16:33:39 2013 |
Sujan | Update | PEM | Guralp2 seismometer installed at North side of POX table | 1)Power to the seismometers were turned down,
2)Guralp2 was moved to North side of POX table
3)Guralp2 was aligned in N-s Direction and leveled before connecting
4)Power to seismometers was turned on once Guralp2 was connected

|
8953
|
Thu Aug 1 16:03:52 2013 |
Steve | Update | General | reopen BS & IOO |
Quote: |
[Koji, Manasa]
We missed to check that we had the green transmitted to the PSL after flipping the SRC and PRC folding mirrors.
There is no green transmission reaching the PSL even after locking the arms to green.
We should fix this tomorrow. The BS heavy door should come off.
Steve! Do not start pump down tomorrow !
|
We closed the chambers last night with heavy doors and reopened it today. |
8952
|
Thu Aug 1 15:28:44 2013 |
gautam | Update | CDS | New model for endtable PZTs-problem solved |
Quote: |
I don't know what's going on here (why the channels are white), and I don't yet have a suggestion of where to look to fix it but...
Is there a reason that you're making a new model for this? You could just use and existing model at c1iscex, like the c1scx, and put your stuff in a top-names block. Then you wouldn't have to worry about all of the issues with adding and integrating a new model.
|
Koji just fixed this.
It seems that the new model's channels were not automatically added to the master file in the framebuilder (/opt/rtcds/caltech/c1/target/master). Adding the following two lines to the master file fixed the problem;
/opt/rtcds/caltech/c1/chans/daq/C1ASX.ini
/opt/rtcds/caltech/c1/target/gds/param/tpchn_c1asx.par
The box is now green. It looks like C1ASX.ini is created automatically in /opt/rtcds/caltech/c1/chans but the master file needs to be manually edited. The channels are now showing up on dataviewer etc. I have updated the information on the wiki page.
|
8951
|
Thu Aug 1 15:06:59 2013 |
jamie | Update | CDS | New model for endtable PZTs |
Quote: |
I have made a new model for the endtable PZT servo, and have put it in c1iscex. Model name is c1asx. Yesterday, Koji helped me start the model up. The model seems to be running fine now (there were some problems initially, I will post a more detailed elog about this in a bit) but some channels, which are computer generated, don't seem to exist (they show up as white blocks on the MEDM GDS_TP screen). I am attaching a screenshot of the said screen and the names of the channels. More detailed elog about what was done in making the model to follow.

Channel Names:
C1:DAQ-DC0_C1ASX_STATUS (this is the channel name for the two leftmost white blocks)
C1:DAQ_DC0_C1ASX_CRC_CPS
C1:DAQ-DC0_C1ASX_CRC_SUM
|
I don't know what's going on here (why the channels are white), and I don't yet have a suggestion of where to look to fix it but...
Is there a reason that you're making a new model for this? You could just use and existing model at c1iscex, like the c1scx, and put your stuff in a top-names block. Then you wouldn't have to worry about all of the issues with adding and integrating a new model. |
8950
|
Thu Aug 1 13:09:17 2013 |
gautam | Update | CDS | New model for endtable PZTs-procedure |
These are roughly the steps I followed in setting up the new model for the endtable PZT servo - C1ASX.
Simulink model:
I made a SIMULINK model of the servo, using MATLAB R2013a. The path to the model is /opt/rtcds/caltech/c1/userapps/release/isc/c1/models/c1asx.mdl. I am listing the parameters set on the CDS_PARAMETERS block:
- host = c1iscex
- site = c1
- rate = 16k
- dcuid = 44 (which I chose after making sure that this dcuid was not used on this list which was last updated end Feb 2013)
- specific_cpu = 5 (again chosen after checking the available CPUs in the above list).
- adc_Slave = 1
- shmem_daq = 1
- no_rfm_dma = 1
- biquad = 1
Making, Compiling and Installing the Model:
After saving the model, I ssh-ed into c1iscex and ran the following commands:
rtcds make c1asx - this gave me a whole bunch of errors initially, which I tracked down to a naming problem in some of the from and goto flags: there should not be any spaces.
rtcds install c1asx
rtcds start c1asx - this gave me an error which said something like 'can't start/stop model.' Koji pointed out that given that a new model is being started, there is an additional step involved, which is to add the model name to the rtsystab file (this is located at /diskless/root/etc/rtsystab on framebuilder, and is mirrored in the various computers. It would be advisable to make sure that the changes are mirrored in the corresponding file on the computer in which the new model is being installed).
After adding the model name to the rtsystab file, I tried running rtcds start c1asx again. This time, no errors were output, but the model was not up and running as verified by looking at the C1:ASX_GDS_TP medm screen.
Debugging
Koji suggested making a simple model (1 CDS parameters block, 1 ADC block and 2 filter modules, appropriately terminated) and see if that starts up, which it did. I then tried adding my servo minus the DAC block and recompiled and restarted the model. This too worked fine. I figured that the next logical step would be to add the DAC block to the model, and restart the model. But when I tried this, c1iscex crashed .
Jenne helped in restoring things to a working state (we reverted the c1asx model to just 2 filter modules, and went to the X-end and restarted the computer. This did not work the first time so I went back in and restarted it again, at which point we were able to ssh into c1iscex again and restart the four models running on it).
Since Manasa and Koji were working on getting things set up for the pumpdown,I did not try anything again till later in the evening, when Koji helped in debugging the problem further. In the meantime, at Jenne' suggestion, I made the model once again in MATLAB R2010b. In the evening, when I tried restarting the model, Koji suggested that the DAC channels in c1asx may be used by other models, at which point I realised I had set up excitation points on channels 8 through 15 of the DAC in c1scx (detailed here) in order to test the hardware at 1X9. I removed the excitation points from channels 8-11 of the DAC block in c1scx (these are the ones used in c1asx), and recompiled and restarted c1asx (using the above sequence of commands). I then tried recompiling and starting c1asx once more, and this time, it worked . At least, the GDS_TP screen suggests that the model is running alright, except for the fact that some computer generated channels seem to be missing. This problem is unresolved for now, and probably has something to do with the fact that C1ASX channels do not appear in Dataviewer.
I do not think this has to do with restarting framebuilder (I did the usual telnel fb 8088 followed by shutdown). In any case, I have added the new model to the CDS_FE_STATUS screen, and will continue to debug the same. I have also got a template medm screen (work in progress) which I will elog about soon as I get it done.
Note to self: There are 4 more excitation channels still hooked up to the DAC (channels 12-15) in the c1scx model. I plan to remove these and put them in c1asx.
|
8949
|
Thu Aug 1 12:12:35 2013 |
gautam | Update | CDS | New model for endtable PZTs | I have made a new model for the endtable PZT servo, and have put it in c1iscex. Model name is c1asx. Yesterday, Koji helped me start the model up. The model seems to be running fine now (there were some problems initially, I will post a more detailed elog about this in a bit) but some channels, which are computer generated, don't seem to exist (they show up as white blocks on the MEDM GDS_TP screen). I am attaching a screenshot of the said screen and the names of the channels. More detailed elog about what was done in making the model to follow.

Channel Names:
C1:DAQ-DC0_C1ASX_STATUS (this is the channel name for the two leftmost white blocks)
C1:DAQ_DC0_C1ASX_CRC_CPS
C1:DAQ-DC0_C1ASX_CRC_SUM |
8948
|
Wed Jul 31 21:12:05 2013 |
manasa | Update | General | Pump down called off | [Koji, Manasa]
We missed to check that we had the green transmitted to the PSL after flipping the SRC and PRC folding mirrors.
There is no green transmission reaching the PSL even after locking the arms to green.
We should fix this tomorrow. The BS heavy door should come off.
Steve! Do not start pump down tomorrow ! |
8947
|
Wed Jul 31 17:02:17 2013 |
Alex Cole | Update | Electronics | Preliminary Photodetector Frequency Reponse Measurements | [Eric, Alex]
We used our setup from yesterday (elog #8940) to measure transimpedance measurements for AS55, REFL11, REFL33, and REFL55, using our Newport 1611 FC-AC as reference. We connected the fibers to their respective telescopes such that the beams focused on their photodetectors, using a multimeter to maximize photodetector DC output. Plots are attached. At first glance, the poles seem to be where they're supposed to be.
Note that the procedure used today is similar to what the eventual automated procedure will be. The main differences are (1) The RF Switch will be used rather than manual switching (2) NWAG4395A will be used to collect data rather than netgpibdata (3) Data will be fit using vectfit4.m and compared to some canonical set. |
8946
|
Wed Jul 31 10:27:00 2013 |
Steve | Update | VAC | instrument grade air is discontinoued | We have been using Air Liquide Instrument Grade Air for venting the 40m vacuum envelope. It is no longer available.
The replacement is Alphagas 1, total hydrocarbon <0.1% ppm
LINK TO Alphagaz

Air Liquide > ALPHAGAZ™ Specifications

ALPHAGAZ™ Specifications
ALPHAGAZ™ range of Pure Gases for Analysis: two purity levels tailored to cover all your analytical needs.
Because each analytical application is sensitive to critical impurities that can affect your results, tests in our research center and analyzer manufacturers' studies confirm that the main impurities are moisture, oxygen and hydrocarbons.
Based on these critical impurities, Air Liquide has developed ALPHAGAZ™ brand with a Maximum Impurity Level (H2O, O2, CnHm).
- The Maximum Impurity Level is the same across the whole product range. This constitutes a guarantee of quality with regard to major critical impurities and contamination tracers. ALPHAGAZ™ 2 offers an even greater guarantee in terms of purity (CO, CO2, H2).
- The Minimum Total Purity is a prerequisite for product quality, but we do not stop there: Air Liquide selects and controls its gas sources by defining specifications regarding impurities that are critical for your analysis or application.
ALPHAGAZ™ 1 |
ALPHAGAZ™ 1 |
ALPHAGAZ™ 2 |
ALPHAGAZ™ 2 |
ALPHAGAZ™ AUTO IV |
ALPHAGAZ™ CO2 SFC |
He, N2, H2,
Ar, O2 |
Air |
He, N2, H2,
Ar, O2 |
Air |
Air*, N2, O2 |
CO2 |
Specifications (ppm vol): |
H2O < 3 ppm
O2 < 2 ppm
CnHm < 0.5 ppm |
H2O< 3 ppm
CnHm< 0.1 ppm
CO2 < 1 ppm
CO < 1 ppm |
H2O < 0.5 ppm
O2 < 0.1 ppm
CO2 < 0.1 ppm
CO < 0.1 ppm
CnHm < 0.1 ppm
H2 < 0.1 ppm |
H2O < 0,5 ppm
CnHm < 50 ppb
CO2 < 0,1 ppm
CO < 0,1 ppm
NOx < 10 ppb *
SO2 < 10 ppb* |
CnHm < 0.2 ppm
CO2 < 0.5 ppm
CO < 0.5 ppm
NO/NOx < 0.1 ppm |
H2O < 5 ppm
O2 < 2 ppm
CnHm < 2 ppm
NVOC* < 2 ppm
CO < 5 ppm
NO + NO2 < 2 ppm
S Total < 1 ppm
H2 < 0.5 ppm |
|
20 ± 1 % O2 (Europe)
21 ± 1% O2 (North America) |
|
20,9 ± 1 % O2
*: Detection limit |
*20 ± 1 % O2 |
*NVOC: Non Volatile Organic Compounds |
Minimum Total Purity |
|
8945
|
Wed Jul 31 09:51:43 2013 |
Steve | Update | safety | Sujan receives safety training | Sujan got 40m specific basic safety training this morning. |
8944
|
Wed Jul 31 00:59:58 2013 |
manasa | Update | General | ITMX chamber vented/Closing plan | My Plan for closing:
Today: I will work on the ITMX oplev situation today and go through the vent close-up checklist as far as I can get.
[Alex, Sujan, Manasa]
The ITMX oplev steering mirrors were laid out such that they were out of the way of the BS oplev leakage. But the halo associated with the He-Ne laser does exist even now. I conclude that this is something that can be dealt with after we pump down as well. So I did not change the ITMX oplev optics on the POX table.
BS, PRM, ITMY and SRM oplevs were aligned and centered.
We want to do IPANG and IPPOS alignment when the IFO is aligned satisfactorily and right before we put the heavy doors.
The arms were aligned and ASS'd before I went in to fix the oplevs. I haven't done anything but deal with the oplevs tonight. So I am being lazy by assuming the alignment is still good and calling it a night.
Tomorrow: We will do the final alignment check for the arms, PRC and SRC with the light doors on. Check IPANG and IPPOS. The access connector and heavy doors should go in place late afternoon.
Thursday: We will start pumping down early in the morning on Thursday. |
8943
|
Tue Jul 30 19:44:05 2013 |
gautam | Configuration | endtable upgrade | Second mirror glued to PZT and mounted |
I have glued a fourth mirror to a PZT (using superglue) and inserted it into a modified mount. This is to be used together with the 1-inch Laseroptik mirror I had glued a couple of weeks back at the Y-endtable. I will be calibrating both these mirrors tonight such that these are ready to put in as soon as we are pumped down.
The mirror was one of those removed from the X-endtable during the switch of the steering mirrors. It is a CVI 2-inch mirror, with HR and AR coatings for 532 nm. |
8942
|
Tue Jul 30 19:40:47 2013 |
gautam | Configuration | endtable upgrade | DAC-PZT Driver Board Output Signal Chain Tested |
[Alex, Gautam]
The signal chain from the DAC output to the output of the PZT driver board (including the HV supply) has been verified.
I had installed the two boards in the eurocrate yesterday and laid out the cables from 1X9 to the endtable. The output of the AI board had been verified using the monitor port on the front panel, but the output from the PZT driver board was yet to be checked because I had not connected the HV supply yesterday.
When I tried this initially today, I was not getting the expected output from the monitor channels on the front panel of the PZT driver board, even though the board was verified to be working. Alex helped debug the problem, which was identified as the -15V supply voltage not making it onto the board.
I changed the slot the board was sitting in, and used a long screw to bolt the board to the crate. Both the AI board and the PZT driver board seem to be slightly odd-sized, and hence, will not work unless firmly pushed into the eurocrate and bolted down. This would be the first thing to check if a problem is detected with this system.
In any case, I have bolted both boards to the eurocrate, and the output from the PZT driver board is as expected when I sent a 10Vp sine wave out from the DAC. I think the cables can now be hooked up to the PZTs once we are pumped down. |
8941
|
Tue Jul 30 18:56:31 2013 |
manasa | Update | General | ITMX chamber vented/Closing plan | [Koji,Manasa]
We removed the ITMX heavy door to fix the oplev situation.
My Plan for closing:
Today: I will work on the ITMX oplev situation today and go through the vent close-up checklist as far as I can get.
Tomorrow: We will do the final alignment check tomorrow with the light doors. The access connector and heavy doors should go in place late afternoon.
Thursday: We will start pumping down early in the morning on Thursday. |
8940
|
Tue Jul 30 16:21:46 2013 |
Alex Cole | Update | Electronics | Photodetector Input Modulation | [Eric, Alex]
We successfully used our system to modulate the input to a single photodetector. The RF Out of the network analyzer went to the Mod In of our laser, which was operating at 98 mA. The laser's output was sent to our 1x16 optical splitter. This provided input signals for both our reference detector and AS55. Our reference detector's output was sent to the network analyzer's R input, while the AS55's output was sent to the network analyzer's A input.
We still need to work out the specifics of how the modulation works. Specifically, we want to look at the amplitude of the network analyzer's output. Additionally, we may have been saturating our reference detector, causing noise problems. |
8939
|
Tue Jul 30 15:38:24 2013 |
manasa | Update | General | ITMX oplev |
Quote: |
I wonder what optics is causing the halo on the oplev beam.
It this comes from any uncoated lens (or similar) it should be identified.
|
We identified this to be coming from the uncoated concave lens that we have right after the He-Ne laser which should also be replaced in addition to the other problems with the oplev. |
8938
|
Tue Jul 30 13:41:35 2013 |
Koji | Update | General | ITMX oplev | I wonder what optics is causing the halo on the oplev beam.
It this comes from any uncoated lens (or similar) it should be identified. |
8937
|
Tue Jul 30 11:30:55 2013 |
manasa | Update | General | ITMX oplev |
Quote: |
[Koji, Manasa, Sujan]
The ITMX oplev was also checked from the windows without removing the door.
The beam is actually hitting the right positions of the steering mirror and the test mass
although there are lots of stray beams because of the BS oplev and some halo assciated with the ITMX He-Ne laser(!?).
|
[Steve, Gautam, Manasa]
While we checked the ITMX oplev situation yesterday, we found that the beam hitting the ITM and the in-vacuum steering mirrors had a halo around them. We used the set of irides in the path of the ingoing beam and cut the stray light around the beam. This reduced the intensity of the halo around the mirrors. We noticed that the halo accounted for 2000 counts of 6500 at the oplev QPD. We tried changing the laser and this did not make the situation any better.
Also there are a couple or more strong stray beams from the BS oplev.
Thoughts:
I suspect that the BS oplev leakage is messing with the ITMX oplev. Why?? We have been seeing the breathing of transmitted beam from the X arm cavity and the shadow sensor readings have been bigger than usual since the last vent (when we changed the BS oplev path). Also, the ingoing oplev beam is close to clipping at the PR2 stack.
I think it would be best to open the ITMX chamber and modify the in-vac steering mirror layout of the oplev. |
8936
|
Tue Jul 30 00:21:15 2013 |
Koji | Update | General | Vent progress | [Koji, Manasa, Sujan]
Tomorrow we'll make final checks of the optics inside the chamber.
Then we will pump down the chamber.
- AS beam steering
The AS beam was aligned on the steering mirror (Piezo Jenna i.e. PZT4).
At this point we already found the beam on the AS CCD.
After checking the spots in the OMC chamber, we determined the beam is not clipped inside.
On the AS table, there is a weird thick lens with short focal length (like f=37.5mm), this actually clipped the beam significantly
[I take the blame for it from the last vent. The lens was installed thinking it would be temporary which it was not. -Manasa].
We removed this lens and put the CCD camera back so that the spot on the camera became smaller.
- SRM debiasing
First of all, the bias of the SRM was moved to zero.
Then the tower was rotated so that the beam is aligned in yaw.
After fine alignment of the SRM, the yaw slider was changed from -5.8 to 1.2.
The OSEMs on the SRM suspension was also checked. The OSEM insertions were
adjusted so that they show the half values of the open calibration
(the values can be found on the 40m wiki)
SRM / ITMY / BS / PRM oplev paths in the chamber were checked.
The beams were aligned on the steering mirrors and the QPDs on the optical table.
The ITMX oplev was also checked from the windows without removing the door.
The beam is actually hitting the right positions of the steering mirror and the test mass
although there are lots of stray beams because of the BS oplev and some halo assciated with the ITMX He-Ne laser(!?).
- Checking the alignment
After we closed the light doors, we aligned the MC, the arms and the recycling mirrors.
Now the everything is aligned. |
8935
|
Mon Jul 29 21:57:45 2013 |
gautam | Configuration | endtable upgrade | Hardware installed at 1X9 | The following hardware has been installed on rack 1X9;
- KEPCO high voltage power supply (kept in a plastic box at the bottom of the rack, with the 3m SMA cable carrying 100V running along the inside side wall of the rack). The HV supply has not been connected to the driver board yet.
- AI board D000186 installed in top eurocrate. The board does not seem to fit snugly into the slot, so I used a longish screw to bolt the front panel to the eurocrate.
- PZT driver board D980323 installed in top eurocrate adjacent to the AI board.
- Six 11m SMB-LEMO cables have been laid out from 1X9 to the endtable. I have connected these to the PZT driver board, but the other end (to the PZTs) is left unconnected for now. They have been routed through the top of the rack, and along the cable tray to the endtable. All the cables have been labelled at both ends.
I have also verified that the AI board is functional in the eurocrate by using the LEMO monitoring points on the front panel.
The driver boards remain to be verified, but this cannot be done until we connect the HV supply to the board.
|
8934
|
Mon Jul 29 16:15:13 2013 |
Koji | Update | VAC | AC removed | Prior to the access connector removal, Manasa and I aligned the IFO mirrors.
The arms were locked and aligned by ASS. |
8933
|
Mon Jul 29 16:04:40 2013 |
Steve | Update | VAC | AC removed | Manasa, Eric, Evan, Koji and Steve,
Access connector removed in order to complete alignment. Light aluminum with acetate windows AC installed.
|
8932
|
Mon Jul 29 13:39:25 2013 |
gautam | Configuration | endtable upgrade | PZT Driver Board-further changes |
I have updated the schematic of the D980323 PZT driver boards to reflect the changes made. The following changes were made (highlighted in red on the schematic):
- Gain of all four HV amplifier stages changed from ~15 to ~5 by swapping 158k resistors R43, R44, R69 and R70 for 51k resistors.
- Electrolytic 10 uF capacitors C11, C12, C29 and C31 swapped for 470pF, 500V mica capacitors.
- Fixed resistor in voltage divider (R35, R40, R59 and R64) replaced with 0 ohm resistors so as to be able to apply a bias of -10V to the HV amplifier
- The DC-DC Series components, which I think were originally meant to provide the 100V DC voltage, have been removed.
- The path between the point at which +100V DC is delivered and jumpers J3 and J6 has been shorted (bypassing R71 and R11 for J3, R73 and R12 for J6).
- Tantalum capacitors C38 and C39 have been replaced with electrolytic capacitors (47 uF, 25V). One of the original tantalum capacitors had burned out when I tried installing the board in the eurocrate, shorting out -15V to ground. At Koji's suggestion, I made this switch. The AD797s do not seem to be oscillating after the switch.
I have also changed the routing of the 100V from the HV power supply onto the board, it is now done using an SMA T-connector and two short lengths of RG58 cable with SMA connectors crimped on.
The boards are functional (output swings between 0 and 100V as verified with a multimeter for input voltages in the range -10V to +10V applied using a function generator.
Revised schematics:


|
8931
|
Sun Jul 28 20:13:41 2013 |
Annalisa | Update | Green Locking | Y Phase tracker calibration | I did a calibration measurement for the Y part of the BeatBox using a Marconi. This is in order to get a more accurate calibration for the arm cavity scan measurement.
The calibration factor I found is:
C1:ALS-BEATX_FINE_PHASE_OUT 50.801 +/- 0.009 deg/MHz
Procedure
During my cavity scan measurement, I had recorded the beat frequency and amplitude from the Spectrum Analyzer at each zero crossing.
I connected the Marconi to the RF in of the Y part of the BeatBox, and I set the Marconi carrier frequency at one of this zero-crossing frequency that I had recorded, while I set the amplitude in way to have on the spectrum analyzer the same beat amplitude that I read during the measurements or, equivalently, in order to have C1:ALS_BEATY_FINE_Q of the order of 1200 (which is the same value I had during my measurements).
I started with
- Carrier frequency = 80.2 MHz
- Amplitude = -3dBm
Then I monitored the C1:ALS_BEATY_FINE_I on the oscilloscope and I adjusted the carrier frequency so that I had zero signal on the oscilloscope. Eventually the frequency corresponding to the zero crossing was 79.989 MHz.
I resetted the phase (clear history in the BEATY_FINE_PHASE panel) and I started changing the frequency by steps of 0.2 MHz, and I spanned about 70 MHz (from 32 to 102 MHz).
Resutls
The calibration coefficient I found is not so different from the one that Yuta measured (elog 8199).
Here are the fit parameters:
y = a + bx
a = -4239.7 +/- 0.6 deg
b = 50.801 +/- 0.009 deg/MHz |
8930
|
Sun Jul 28 19:39:04 2013 |
Annalisa | Update | endtable upgrade | Picture | Yend table picture updated on the wiki page |
8929
|
Fri Jul 26 22:45:14 2013 |
manasa | Update | General | Vent progress | [Koji, Manasa]
The arms were brought back to resonance after yesterday's vent work.
DCPD gains were changed (TRX gain from -0.002 to -0.04; TRY gain changed from 0.002 to 0.04) to lock the arm with low power.
X arm and Y arm were locked for IR.
We then decided to align IPANG because the input pointing seemed good - the right time to fix IPANG. The IPANG spot at the last in-vac steering mirror was aligned to be an inch low in pitch.
We then aligned SRC. SR2 and SR3 were adjusted to center the beam at SRM. SRM was aligned to the retro-reflected beam until we saw flashes. But this position of SRM is not at the good OSEM range. We should correct the SRM suspensions. We postponed this activity for next week and went ahead to look at the status of the AS beam.
Looking through the window and using IR viewer were not convincing enough. We will have to get the access connector out on Monday to do the AS alignment.
We then closed the chamber with light doors and locked the arm cavities again. We brought SRC and PRC to resonance and saw strong flashes at the BS_PRM camera.
To do:
SRM need to be moved in order to remove the big bias in yaw
SRM OSEMS need to be adjusted
Access connector should come out
AS needs to be aligned satisfactorily
ITMX oplev steering mirrors in vacuum should be checked.
|
8928
|
Fri Jul 26 22:19:24 2013 |
Charles | Update | ISS | ISS - Full Schematic - Updated |
Quote: |
I have made significant changes to the ISS schematic, mostly in the form of adding necessary subsystems.
Some changes I have made:
- Added a front page with sheet symbols that are representations of the other schematic sheets.
- Added an 'Excitation' subsystem for use in determining the closed-loop transfer function
- Added an instrumentation amplifier (with ADA4004s at Rana's recent suggestion) to handle the differential input from the PD
- Included a switchable inverting amplifier (Gain of 1 or -1) to ensure we have the correct polarity
- Made it so the first filtering stage is immediately active when the ISS loop is closed
- Added LP filters with large time constants to buffer/delay trigger signals
- Added test points all over the board
- Refined a few buffer amplifiers
On the front page, all inputs and outputs are currently BNC ports, although this is most likely not the final design that will be used. For instance, the ports ENABLE, INPUT GND and INVERT are supposed to be logic inputs for a MAX333a switch. These will most likely be front panel switches that either connect the switch's logic pin to GND (Logic 0) or something like a +5 V supply (Logic 1).
I also have not included power regulation for my board although I have some of the actual D1000217 Chasis Power Regulator boards and I'll incorporate those in my design soon.
|
More changes that I've made:
- Added daughter boards for power regulation. Currently I have ±24V going into two boards, with ±15V coming out of one and ±5V coming out of the other. Again, these are based off of LIGO-D1000217
- Added an optional Dewhitening filter (with p=1Hz and z=100Hz, although these can easily be changed) to accommodate any PD's that have whitening
- Added a bypass to allow the boosts (stages 2 and 3 of the filtering servo) to be enabled/disabled by a front panel switch
- I also put in jumpers that can be used to provide Logic 1 (boost enabled) to both Boost 1 and Boost 2 without depending on the internal RMS detection/triggering
- Changed the input grounding switch so that it's set up correctly. Before, it was taking the PD signal and sending it to GND, not actually grounding the input to the rest of the ISS
|
8927
|
Fri Jul 26 14:39:08 2013 |
Charles | Update | ISS | Power Regulation for ISS Board | I constructed a regulator board that can take ±24 V and supply a regulated ±15 V or ±5 V. I followed the schematics from LIGO-D1000217-v1.
I was going to make 2 boards, one for ±15 V and one for ±5, but Chub just gave me a second assembled board when I asked him for the parts to construct it 
|
8926
|
Fri Jul 26 13:39:30 2013 |
Koji | Update | ASC | POP QPD analysis | It was not actually easy to see from the entry what signal was taken in what condition but from the shape of the spectra
I had the impression that the ASC & OPLEV signals were measured under the presence of the ASC control.
That is (moderately to say) tricky as the ASC control imprints the angular noise
from unkown mirror on the PRM, and then the oplev observes it. The original stability of the oplev is
obscured by the injection from the servo and the fair comparison of the stability is almost impossible.
So the true comparison between the ASC and oplev signals should be done without the control loop.
http://nodus.ligo.caltech.edu:8080/40m/8532
http://nodus.ligo.caltech.edu:8080/40m/8535
We can recover the free running spectrum of the ASC signals by compensating the loop transfer functions
because the ASC signals are the in-loop error signals. The oplev signals should be measured without
the ASC loop engaged. |
8925
|
Thu Jul 25 14:45:06 2013 |
rana | Update | PEM | Guralp specgram with ligoDV and NDS2 | Once you install a matlab newer than 2012a, you can install ligoDV as a matlab app and get the NDS2 client software for free. So you can easily get the 40m data from the outside world now and do the analysis on your own computer rather than login through nodus. |
8924
|
Thu Jul 25 14:02:53 2013 |
Jenne | Update | SUS | SR3, PR3 flipped | Yesterday afternoon, I went back into the BS chamber, and flipped both PR3 and SR3. Now all of the recycling cavity folding mirrors have been flipped.
For PR3, I followed the same procedure as SR2, setting a reference position, removing the optic, flipping it, etc. When I put it back in, I realized that since this has a 41 degree angle of incidence, the beam going to the BS had translated north by ~1cm. After some fiddling, Koji pointed out that the 2 degree wedge probably had a more significant effect than just the HR surface having moved back a small amount. Anyhow, we adjusted PR3 such that we were going through the BS aperture, as well as the ITMY aperture.
During the flip of PR3, Annalisa and I noticed that the arrow on the barrel of the LaserOptik mirrors also indicates the thickest part of the wedge. This is opposite of our SOS optics, where the arrow's position on the barrel indicates the thinnest part of the wedge. For both PR3 and SR3, I kept the arrow on the same side of the optic as it was originally.
I then flipped SR3, following again the same procedure. PR3 I had done a tiny bit of pitch rebalancing, although I think it was unneccessary, since it is within what we can do with the poking/hysterisis method. SR3 I did not do any pitch rebalancing. With PR3 aligned at least to the ITM, Koji and I aligned SR3 and SR2 so that the AS beam was hitting the center of all the SRC optics. We also adjusted the steering mirrors after the SRM to get the beam centered on PZT3, the last optic on the BS table, which launches the beam over to the OMC chamber. We scanned around a bit by turning the PZT's knobs, but we were unable to see the AS beam on the camera.
|
8923
|
Thu Jul 25 13:54:35 2013 |
manasa | Update | General | PR3 clamped and Y arm is back flashing | [Jenne, Annalisa, Manasa]
After yesterday's flipping of PR3, we lost our input pointing. Koji spent a few hours last night but couldn't restore the Y arm. I did my set of trials this morning which also didn't help.
So Jenne and I went ahead and requested Steve to get the ETMY door off.
We set the tiptilts TT1 and TT2 to the slider values from yesterday and started aligning the PR3 to hit the center of ITMY.
When we were hitting close to the center of ITMY, we decide to use the tip-tilts because the movement of PR3 was coarse at this point.
We used TT1 to get the beam to the center of ITMY and TT2 to get the beam at the center of ETMY. We did this iteratively until we were at the center of both the ITMY and ETMY.
We then went to fix IPANG.
The IPANG steering mirror on the BS table was steered to hit the center of the steering mirrors at the ETMY table. We aligned the beam to the IPANG QPD on the green endtable. The steering mirror on the BS table was then steered to misalign the beam in pitch by an inch at the last IPANG steering mirror. This should fix the IPANG clipping we have everytime we pump down.
We closed the chambers with light doors and saw IR flashing in the arm cavity. Koji is now trying to lock the cavity with IR. |
8922
|
Thu Jul 25 12:53:45 2013 |
Charles | Update | ISS | Comparator + Triggering Prototype | I realized I totally forgot to post this last week, but I prototyped the comparator and boost triggering portion of the ISS, at least in part. Below is a schematic that shows the prototype circuit I made. Note that it includes ports for the oscilloscope channels that appear in the second image included. Essentially, I was able to verify that the output from the LT1016, as it's currently constructed in the ISS schematic, would be sufficient logic to switch the MAX333a.

Below, we can first see that the comparator is switching its output as desired. When the DC level of the input drops below a certain threshold (~1.6 V) the output of the comparator switches on to ~4 V. When the DC level of the input goes back up above the upper threshold (~3.2 V), the comparator switches off to ~0.3 V. The exact values of the threshold voltages can be determined/tuned at a later date, but this is the basic behavior that the comparator circuit will have.
To detect whether or not the MAX333a was switching properly, I connected the common terminal of one of the switches to a +5 V supply, and looked at the voltage coming off both the 'open' and 'closed' terminals of said SPDT switch. We can see that with Logic 0 (comparator output ~0.3 V) Channel 4 exhibits a ~5 V signal, just as we would expect from the above schematic. With Logic 1 (comparator output ~4 V), Channel 3 exhibits the characteristic 5 V signal.

|
8921
|
Thu Jul 25 02:53:00 2013 |
Koji | Update | General | IFO Alignment after TT flipping - no progress | There was no progress tonight after Jenne left.
I could not find any reasonable fringes of the IFO after 3 hours of optics jiggling.
* I jiggled TT1 and TT2. The slider has not been restored.
We should probably look at the value in the day time and revert them.
(Still this does not ensure the recovery of the previous pointing because of the hysteresis)
* The arms are still aligned for the green.
It's not TEM00 any more because of the vent/drift but the fringe is visible (i.e. eigenaxis is on the mirror)
* As we touched PR3, the input pointing is totally misaligned.
To Do / Plan
* We need to find the resonance of the yarm by the input TTs. Once the resonance is found, we will align the PRM.
* Move the BS to find the xarm resonance.
* Finally align SRM
* It was not possible to find the resonance of the yarm without going into the chamber. Definitely we can find the spot on the ITMY by a card, but we are not sure the beam can hit the ETMY. And the baffles makes the work difficult.
* One possibility is to align the input beam so that the ITMY beam is retroreflected to the PRM. I tried it but the beam was not visible form the camera. |
8920
|
Wed Jul 24 22:58:03 2013 |
Charles | Update | ISS | ISS - Full Schematic - Updated | I have made significant changes to the ISS schematic, mostly in the form of adding necessary subsystems.
Some changes I have made:
- Added a front page with sheet symbols that are representations of the other schematic sheets.
- Added an 'Excitation' subsystem for use in determining the closed-loop transfer function
- Added an instrumentation amplifier (with ADA4004s at Rana's recent suggestion) to handle the differential input from the PD
- Included a switchable inverting amplifier (Gain of 1 or -1) to ensure we have the correct polarity
- Made it so the first filtering stage is immediately active when the ISS loop is closed
- Added LP filters with large time constants to buffer/delay trigger signals
- Added test points all over the board
- Refined a few buffer amplifiers
On the front page, all inputs and outputs are currently BNC ports, although this is most likely not the final design that will be used. For instance, the ports ENABLE, INPUT GND and INVERT are supposed to be logic inputs for a MAX333a switch. These will most likely be front panel switches that either connect the switch's logic pin to GND (Logic 0) or something like a +5 V supply (Logic 1).
I also have not included power regulation for my board although I have some of the actual D1000217 Chasis Power Regulator boards and I'll incorporate those in my design soon. |
8919
|
Wed Jul 24 19:21:56 2013 |
Jamie | HowTo | SUS | SUS MEDM screen modernization | I started poking around at what we want for new SUS MEDM screens. Rana and I decided we'd start with the ASC TIPTILT screens:

It's missing some things (like SIDE OSEMS) but it should provide a good starting point.
I copied the entire <userapps>/asc/common/medm/asctt directory to a new directory in our sus area:
controls@rossa:/opt/rtcds/userapps/release 0$ cp -a asc/common/medm/asctt sus/c1/medm/new
I then removed all the useless file name prefixes. We still need to go through and sed out all the ASC stuff in the MEDM files themselves.
It makes heavy use of macro substitution, which is good (it's what we're using now). So once we clean up all the channel names, we should just be able to swap out the pointers in our overview screens to the new screens (or rename things). In the mean time, during development, you can run:
controls@rossa:/opt/rtcds/userapps/release 0$ medm -x -macro "IFO=C1,ifo=c1,OPTIC=ITMX" sus/c1/medm/new/OVERVIEW.adl
|
|